How many liters of a 30% alcohol solution must be mixed with 90 liters of a 60% solution to get 40% solution?
Answer by checkley77(12844)

.30X+.60*90=.40(X+90)
.30X+54=.40X+36
.30X-,40X=36-54
.10X=-18
X=18/.10
X=180 LITERS OF 30% IS USED.
PROOF:
.30*180+.60*90=.40(180+90)
54+54=.40*270
108=108